Suppose the price of widgets increases from $15 to $20 and the quantity demanded decreases from 100 to 50 units. Calculate the elasticity of demand using the midpoint method. Round your answer to the nearest 2 digits and do NOT take the absolute value.
Answer #1

Elasticity of demand= 6 -Q, ) +P, ) (P₂ - P. (Q2 + 6) (50-100 ) (20+15) (20015) (50t1oo) - 50 x 35 150 = -2-33